Price per kg Quantity demanded

R30 1000

R32 850

Using the midpoint method calculate the price elasticity of demand (using 2 decimal places) for the R30 to R32 price range. The price elasticity of demand for tomatoes is

1
Expert's answer
2021-07-29T14:09:04-0400

The price elasticity of demand for tomatoes is:

"Ed = \frac{850 - 1000} {32 - 30} ×\frac{32+30} {850+1000} = -2.51."

So, the demand is elastic.